Skip to end of metadata
Go to start of metadata

Last (part) day in the field – I am writing this from Lisbon...

All but 2 masts that have been constructed are now powered up and collecting data.  There are 7 masts remaining to be prepped in the ops center.  There are 2? masts that have DTU sensors but have not yet been configured to send their data to our system (easy for them to do once they are on the network).  We also have a handful of sensor issues (mostly RMYoung sonics) that are reporting continuously bad values.  Also, standard maintenance, such as cleaning the radiometers, has not been done (though the rain is performing much of this task!).  We had staged equipment at the two remaining constructed towers, but pulled this equipment off again for safekeeping until we next arrive.

Much of the network is functioning.  HOWEVER, the primary mast that we have set up to get data out of the valley (rne07) failed to pair with the ops center.  Thus we are receiving less than half of the masts that are working.  (rne07 did power up after the quick work by the electricians to resolve a power issue at this site this morning.)  The best guess is that the antenna is not pointed correctly, but time didn't allow us to fix it.  Unfortunately, this means that ISS also is unable to connect with the network, so Gary reinstalled a cell modem at that site.

We also consolidated our mess at the ops center.

The German/Danish group also is leaving today (having placed many of the lidars), so no one will be on site for a while.

 

 

 

 

  • No labels